A partial solution to the equation 5(2x + 3) = 85 is given below. Choose the step that is missing from the solution.
5(2x + 3) = 85

(missing step)
10x = 70
x = 7

Answer by sudhanshu_kmr(1152)


5(2x + 3) = 85
=> 10x + 15 = 85
=> 10x = 85-15

=> 10x = 70
=> x = 7